Consider the reaction A + B ---> products


From the following data obtained at a certain temperature, determine the rate law, the order of the reaction, and calculate the rate constant k.


Experiment 1: [A] = 1.50 M; [B] = 1.50 M; Initial Rate = 3.20 x 10-1 M/s


Experiment 2: [A] = 1.50 M; [B] = 2.50 M; Initial Rate = 3.20 x 10-1 M/s


Experiment 3: [A] = 3.00 M; [B] = 1.50 M; Initial Rate = 6.40 x 10-1 M/s


1
Expert's answer
2020-11-18T13:45:16-0500

"Rate = k[A]^a[B]^b"

"\frac{Rate_1}{Rate_2} = (\frac{[A_2]}{[A_1]})^a \times (\frac{[B_2]}{[B_1]})^b"

"\frac{3.2 \times 10^{-1}}{3.2 \times 10^{-1}} = (\frac{[1.5]}{[1.5]})^a \times (\frac{[2.5]}{[1.5]})^b"

"1 = 1^a \times 1.67^b"

b =0

"\frac{Rate_3}{Rate_1} = (\frac{[A_3]}{[A_1]})^a \times (\frac{[B_3]}{[B_1]})^b"

"\frac{6.4 \times 10^{-1}}{3.2 \times 10^{-1}} = (\frac{[3.0]}{[1.5]})^a \times (\frac{[1.5]}{[1.5]})^b"

"2 = 2^a \times 1^b"

a = 1

Rate = k[A]

Overall order of reaction = a + b = 1 + 0 = 1

Reaction is the first order.

Substituting values from Exp. 1 into the rate low:

"3.2 \times 10^{-1} = k \times 1.5"

"k = 0.213 \;s^{-1} = 2.13 \times 10^{-1} \;s^{-1}"